# Uncertainty Quantification of Pre-Trained and Fine-Tuned Surrogate Models using Conformal Prediction

### v1

@@ Abstract
Data-driven surrogate models have shown immense potential as quick, inexpensive approximations to complex numerical and experimental modelling tasks. However, most surrogate models characterising physical systems do not quantify their uncertainty, rendering their predictions unreliable, and needing further validation. Though Bayesian approximations offer some solace in estimating the error associated with these models, they cannot provide they cannot provide guarantees, and the quality of their inferences depends on the availability of prior information and good approximations to posteriors for complex problems. This is particularly pertinent to multi-variable or spatio-temporal problems. Our work constructs and formalises a conformal prediction framework that satisfies marginal coverage for spatio-temporal predictions in a model-agnostic manner, requiring near-zero computational costs. The paper provides an extensive empirical study of the application of the framework to ascertain valid error bars that provide guaranteed coverage across the surrogate model’s domain of operation. The application scope of our work extends across a large range of spatio-temporal models, ranging from solving partial differential equations to weather forecasting. Through the applications, the paper looks at providing statistically valid error bars for deterministic models, as well as crafting guarantees to the error bars of probabilistic models. The paper concludes with a viable conformal prediction formalisation that provides guaranteed coverage of the surrogate model, regardless of model architecture, and its training regime and is unbothered by the curse of dimensionality.
Keywords: Surrogate Models, Uncertainty Quantification, Conformal Prediction
@@ 1 Introduction
Partial Differential Equations (PDEs) governing physical processes are solved using complex numerical modelling simulation codes. Though these codes offer us a mathematically rigid method of solving the relevant equations, the obtained solutions are often limited to discretised domains and require computationally expensive iterative solvers like the finite volume schemes and the finite element methods. Such simulation codes have become central to multiple scientific disciplines in biology (Hospital, Adam et al., 2015), engineering (Giudicelli et al., 2024), and climate science (Danabasoglu et al., 2020) (Lavin et al., 2021), but they possess difficulty in deploying for rapid iterative modelling, often required while exploring a design space of solutions. Machine learning offers an alternative data-driven route for obtaining quick, inexpensive approximations to numerical simulations (Bertone et al., 2019; Karniadakis et al., 2021). Data-driven surrogate models are often built by distilling the spatio-temporal information characterised by the simulations into parameterised machine learning models. Due to the efficiency, cost-effectiveness, and relative accuracy of modelling, neural networks have become ubiquitous within scientific modelling, and they are of primary importance in tackling large-scale PDEs in climate (Lam et al., 2023; Kurth et al., 2023), computational fluid dynamics (Jiang et al., 2020; Pfaff et al., 2021), and nuclear fusion (van de Plassche et al., 2020; Gopakumar and Samaddar, 2020).
@@
As effective as these surrogate models (emulators) are, they remain approximations of the true physical system under study. There are often many layers of approximations as the simulation codes approximate the PDEs and the PDEs themselves are approximations to the true physical system. They often fail to quantify the uncertainty associated with their approximation, i.e. the approximation error of the emulator with that of the numerical code. Irrespective of the domain of operation they were trained on, models often claim to provide confident outputs. This is problematic for two reasons: (a) without an assessment of the confidence of the output, erroneous predictions can lead to severe consequences downstream; (b) The cost of training these models can be prohibitively high, and without appropriate uncertainty quantification the overconfident predictions will have limited utility. Several works have attempted to provide uncertainty estimation for these surrogate models (Geneva and Zabaras, 2020; Alhajeri et al., 2022; Zou et al., 2024; Psaros et al., 2023), but they fail to provide statistical guarantees over the error bars attached to the prediction and to scale to complex scenarios (Abdar et al., 2021). They also require ensemble training Lakshminarayanan et al. (2017), extensive sampling (MacKay, 1992), and perhaps architectural modifications (Gal and Ghahramani, 2016), successively increasing the amount of computational resources required. There remains a concerning question of validating the surrogate model’s outputs for a specific downstream application. Conformal prediction (CP) (Vovk et al., 2005) provides a framework that allows us to compute statistically guaranteed error bars over pre-trained and fine-tuned models i.e. the error bars are calibrated to provide the required coverage. Conformal prediction relies on the calibration of model performance across a specific dataset representative of the desired prediction distribution, and utilising the calibration measures to help provide valid error bars for the model output.
@@
In this paper, we conduct a thorough empirical study on conformal prediction for pre-trained and fine-tuned, neural-network-based surrogate models to equip their predictions with calibrated, guaranteed error bars, even on out-of-training-distribution scenarios. We demonstrate that we can obtain guaranteed coverage across spatio-temporal domains (high-dimensional outputs), irrespective of the choice of model, data and training conditions. Our work provides a rigorous method to check the usefulness of a pre-trained surrogate model, its validity and applicability in an inference/production scenario.
@@
1.1 Pathway to Addressing Climate Change
Machine-learning-based surrogate modelling promises computational efficiency and allows for data-driven discovery at scale. In modelling tasks, such as computational fluid dynamics, nuclear fusion, and weather forecasting, it is increasingly crucial to provide accurate, robust, and timely uncertainty estimates. In such safety-critical systems Knight (2002), it is impervious that we supplement model predictions with calibrated uncertainty estimations. This can lead to improved decision-making processes in downstream tasks. The CP framework demonstrated in this paper pushes the envelope forward in uncertainty quantification for complex scientific modelling at scale, demonstrating industry-level safety-critical applications.
@@ 1.1 Pathway to Addressing Climate Change
Solving Nuclear Fusion at a commercial scale is a key component in decarbonising hard-to-abate industries (Lerede et al., 2023). However, several challenges remain, ranging from confinement(Degrave et al., 2022), and materials (Linke et al., 2019) to mitigating disruptions (Kates-Harbeck et al., 2019). Operating on an accelerated timeline and exponential costs, iterative test-based design becomes an unviable option. Researchers are left to solve these challenges in-silico using simulations. However, within the context of Fusion modelling, at the exascale, simulation costs grow exponentially and need surrogate models capable of quicker, cheaper approximations. Neural-PDE solvers have enabled this, allowing us to solve complex multi-physical systems several orders of magnitude faster than traditional numerical solvers (Pamela et al., 2024; Carey et al., 2024). As further development are done on these models, it is pertinent to be to provide rigorous methods of quantifying the uncertainty associated with these predictive models.
Climate change has led to the observation of extreme weather events across the globe (Ebi et al., 2021). Being able to forecast weather rapidly and cheaply allows us to take proactive rather than reactive measures to these climate change-induced events. Neural weather models learnt from both simulation and observed data have emerged as cheaper and faster alternatives to traditional numerical models. However, they suffer from the limitations found within the training data such as sparse and inaccurate observations as well as the error growth in the numerical models (Sheshadri et al., 2021). Being able to model the uncertainty associated with these neural weather predictions is crucial in an operational context (Bouallègue et al., 2024). Through this work, we propose a model-agnostic method that gives calibrated error bounds for all variables, lead times and spatial locations, while not requiring any modifications to the model albeit with negligible computational costs.
Outline
The paper is outlined as follows: Section 2 introduces the inductive CP framework, spatio-temporal data and its associated exchangeability assumptions. The section also delves into the mathematical framework for the extension of CP for spatio-temoral data. Section 3 contains the experiments of deploying our CP framework across a range of modelling tasks, ranging from PDEs to climate modelling. In Section 4, we discuss the strengths and weaknesses of the framework and conclude our findings in Section 5.
@@ 2 Conformal Prediction
Conformal prediction (Shafer and Vovk, 2008) answers the following question: Given some arbitrary dataset $(X_{1},Y_{1}),(X_{2},Y_{2}),...,(X_{n},Y_{n})$, and some machine learning model $\hat{f}:\mathcal{X}\to\mathcal{Y}$ trained on this dataset, what is the accuracy of $\hat{f}$ at predicting the next true label $Y_{n+1}$ at query point $X_{n+1}$. CP (Vovk et al., 2005) extends the point prediction $\hat{y}$ of $\hat{f}$ to a prediction set $\mathbb{C}^{\alpha}$, which is guaranteed to contain the true label $Y_{n+1}$ with probability
@@
CP is attractive for quantifying errors in machine learning since the inequality in Equation 1 is guaranteed regardless of the selected machine learning model and the training dataset $\left\{(X_{i},Y_{i})\right\}_{i=1}^{n}$, other than that the samples used for the calibration (estimating model performance) are exchangeable (a weaker form of i.i.d.).
There have been several variants of CP since its original proposal by (Vovk et al., 2005). Inductive conformal prediction (Papadopoulos, 2008), which we pursue in this work, splits the data into a training set (the usual training set for the underlying ML model) and a calibration set, used to construct prediction set $\mathbb{C^{\alpha}}$ for each $\alpha$ value. The prediction sets are constructed, such that Equation 1 holds, by comparing the trained ML model to the unseen calibration data using a non-conformity score.
@@
Figure 1: Inductive CP Framework over a deterministic model (see AER in Section 2.2): (1) Perform calibration using a non-conformity metric (L1 error residual (AER) with $\hat{s}$ representing the calibration scores, $y_{c},\tilde{y}_{c}$ the calibration targets and predictions respectively). (2) Estimate the quantile corresponding to the desired coverage from the CDF of the non-conformity scores ($n$ represents the calibration sample size, $(1-\alpha)$ the desired coverage, $F_{\hat{s}}^{-1}$ the quantile function applied over the inverse CDF of non-conformity scores, $\hat{q}$ the quantile matching the desired coverage). (3) Apply the quantile to the model predictions to estimate the prediction sets ($\tilde{y}_{p}$, the model predictions and $\hat{q}$ the upper and lower bars for the predictions).
The inductive CP framework explored further in this paper follows a three-step procedure as outlined in Figure 1. Initially, we perform a calibration of the model outputs, where a non-conformity metric of the model’s performance is evaluated. This is followed by the quantile estimation, which is performed by obtaining the required quantile from the cumulative distribution of the non-conformity scores obtained in the previous calibration step. The final step involves estimating the prediction sets by applying the estimated quantile over the model predictions in the interested prediction regime.
@@ 2.1.2 Overview
The models we consider are those trained to model the evolution of spatio-temporal field variables defined by physical processes, such as those described in numerical PDE modelling, fusion diagnostics and weather forecasting. Each modelling task is perceived as an initial value problem, where the calibration and prediction set are pairs of inputs and outputs characterised by the initial condition and their respective solutions. Within the case of PDE modelling, each datapoint is a single numerical simulation, whereas, for experimental modelling (such as fusion and weather), the datapoint is a single forecast driven by an initial condition.
@@
Within our framework, we perform calibration for each cell individually, estimating the marginal coverage for each cell of the spatio-temporal tensor as demonstrated in Figure 2 i.e., the coverage is guaranteed for each cell taken individually. CP is performed across each spatio-temporal point output by the model, resulting in upper and lower coverage bands for each point. Upon calibrating for the error bars cell-wise, we averaged over each to estimate the coverage across the simulation domain. Within the estimation of the prediction sets, we don’t consider the influence of adjacent field points and implicitly expect the model to extract that within the learning process. We expect the discretised spatio-temporal domain to be the same across the calibration and prediction set.
Within the scope of this paper, a formulation of the CP framework, aiming to provide statistically valid, dimensionality independent, marginal coverage over spatio-temporal prediction models is proposed. We demonstrate the efficacy of our approach through comprehensive and rigorous experimentation across a range of spatio-temporal models from toy PDE problems to complex multi-physics models and state-of-the-art neural weather and Fusion models.
2.1.3 Formal Definition
@@ 2.2 Non-conformity Scores
A non-conformity score can be described as a measure of the model’s performance with respect to the calibration dataset. Thus, the scores are expressed as a function of the trained model, the calibration inputs, and outputs (Angelopoulos and Bates, 2023). Being expressed as a measure of the deviation of the model from ground truth, we focus on three methods of estimating the non-conformity scores for a model:
@@
  • Quantile Regression (CQR): Outlined by (Romano et al., 2019), three models are trained to perform conformal prediction. The lower and upper coverage bands are estimated by models trained to output the $100\times\alpha^{th}$ and $100\times(1-\alpha)^{th}$ percentile, while the median is learned by training the model to output the $50^{th}$ percentile. During training the models are estimated to output the $\alpha^{th}$ percentile by optimising the quantile loss (Koenker, 2005). The non-conformity score measures the distance of the calibration data point from the nearest coverage band: $s(x,y)=\max\{\underline{f}(x)-y,y-\overline{f}(x)\}$. Using the upper and lower quantiles, CQR offers an initial coverage that can be adjusted and calibrated to perform conformal prediction: Using the required quantile ($\hat{q}$), the prediction set is obtained as $\{\underline{f}(x)-\hat{q},\overline{f}(x)+\hat{q}\}$.
@@
  • Absolute Error Residual (AER): Requiring only a single model, trained to approximate the deterministic function itself, this method uses a more straightforward estimate of the model error to calibrate the coverage. Non-conformity scores are estimated by taking the absolute error of the model across the labelled calibration dataset (Lei et al., 2018): $s(x,y)=|y-\tilde{f}(x)|$. Upon deriving the quantile ($\hat{q}$), the prediction set is obtained as : $\{\tilde{f}(x)-\hat{q},\;\tilde{f}(x)+\hat{q}\}$. This method requires no structural changes to the model, and only one deterministic model is required to obtain valid prediction sets. The method uses the distribution of the absolute error residuals (AER) from the model to construct sets with the correct coverage.
@@
  • Standard Deviation (STD): The models are probabilistic, i.e., they predict a distribution rather than a deterministic quantity, where the outputs are characterised by a mean $\mu(x)$ and standard deviation $\sigma(x)$. The non-conformity score, $s(x,y):\frac{|y-\mu(x)|}{\sigma(x)}$, takes into account the uncertainty of the model and conformalises it to provide validity over the predictive uncertainty  [^1] If the choice of the model is not probabilistic in itself, we will have to modify the model with changes to architecture or introduce sampling strategies to output a distribution.. The standard deviation of the output distribution provides estimates of the error bars and, thus is an initial measure of the coverage, which can be adjusted to ensure the correct coverage using conformal prediction. Upon deriving the quantile ($\hat{q}$), the prediction set is obtained as : $\{\mu(x)-\hat{q}\sigma(x),\;\mu(x)+\hat{q}\sigma(x)\}$. To do CP using a probabilistic model, we modify our networks to behave in a probabilistic manner through architectural modifications or changing the training regime.
@@
Figure 3 illustrates how these non-conformity scores can help obtain guaranteed coverage, across all modelling scenarios considered in the paper(more details on them in Section 3). The coverage provided across the non-conformity scores may vary, but the CP formulation ensures the validity of the error bars. Irrespective of our choice of the non-conformity score, we obtain near-to-ideal coverage for the prediction following the inequality found within Equation 1. Since coverage is guaranteed in the framework, the choice of the non-conformity score is often undertaken considering other aspects such as architectural modifications, calibration time, and cost of obtaining additional data.

### v2

# Uncertainty Quantification of Surrogate Models using Conformal Prediction

@@ Abstract
Data-driven surrogate models have shown immense potential as quick, inexpensive approximations to complex numerical and experimental modelling tasks. However, most surrogate models of physical systems do not quantify their uncertainty, rendering their predictions unreliable, requiring further validation. Though Bayesian approximations offer some solace in estimating the error associated with these models, they cannot provide guarantees, and the quality of their inferences depends on the availability of prior information and good approximations to posteriors for complex problems. This is particularly pertinent to multi-variable or spatio-temporal problems. Our work constructs and formalises a conformal prediction framework that satisfies marginal coverage for spatio-temporal predictions in a model-agnostic manner, requiring near-zero computational costs. We provide an extensive empirical study of the application of the framework to ascertain valid error bars that provide guaranteed coverage across the surrogate model’s domain of operation. The application scope of our work extends across a large range of spatio-temporal models, from solving partial differential equations to weather forecasting. Through the applications, the paper looks at providing statistically valid error bars for deterministic models, as well as crafting guarantees to the error bars of probabilistic models. Our conformal prediction formalisation provides guaranteed coverage of the surrogate model, regardless of model architecture, and its training regime and is unbothered by the curse of dimensionality.
Keywords: Surrogate Models, Uncertainty Quantification, Conformal Prediction, Neural-PDE, Neural-Weather
@@ 1 Introduction
Partial Differential Equations (PDEs) governing physical processes are solved using complex numerical modelling simulation codes. Though these codes offer a mathematically-rigorous method of solving the relevant equations, the solutions are often limited to discretised domains and require computationally-expensive iterative solvers like finite-volume schemes and finite-element methods. Such simulation codes have become central to multiple scientific disciplines in biology (Hospital, Adam et al., 2015), engineering (Giudicelli et al., 2024), and climate science (Danabasoglu et al., 2020) (Lavin et al., 2021), but they are difficult to deploy for rapid, iterative modelling, which is often required while exploring a design space of solutions. Machine learning offers an alternative data-driven route for obtaining quick, inexpensive approximations to numerical simulations (Bertone et al., 2019; Karniadakis et al., 2021). Data-driven surrogate models are often built by distilling the spatio-temporal information characterised by the simulations into parameterised machine learning models. Due to the efficiency, cost-effectiveness, and relative accuracy of modelling, neural networks have become ubiquitous within scientific modelling, and they are of primary importance in tackling large-scale PDEs in climate (Lam et al., 2023; Kurth et al., 2023), computational fluid dynamics (Jiang et al., 2020; Pfaff et al., 2021), and nuclear fusion (van de Plassche et al., 2020; Gopakumar and Samaddar, 2020).
@@
As effective as these surrogate models (emulators) are, they remain approximations of the true physical system under study. There are often many layers of approximations as the simulation codes approximate the PDEs and the PDEs themselves are approximations to the true physical system. They often fail to quantify the uncertainty associated with their approximation, i.e. the approximation error of the emulator with that of the numerical code. Irrespective of the domain of operation they were trained on, models often claim confident outputs. This is problematic for two reasons: (a) without an assessment of the confidence of the output, erroneous predictions can lead to severe consequences downstream; (b) The cost of training these models can be prohibitively high, and without appropriate uncertainty quantification the overconfident predictions have limited utility in practice. Several works have attempted to provide uncertainty estimation for these surrogate models (Geneva and Zabaras, 2020; Alhajeri et al., 2022; Zou et al., 2024; Psaros et al., 2023), but they fail to provide statistical guarantees over the error bars attached to the prediction and to scale to complex scenarios (Abdar et al., 2021). They also require ensemble training Lakshminarayanan et al. (2017), extensive sampling (MacKay, 1992), or architectural modifications (Gal and Ghahramani, 2016), successively increasing the amount of computational resources required. There remains a concerning question of validating the surrogate model’s outputs for a specific downstream application.
Conformal prediction (CP) (Vovk et al., 2005) provides a framework that allows us to compute statistically guaranteed error bars over pre-trained and fine-tuned models, i.e. the error bars are calibrated to provide the required coverage. Conformal prediction relies on the calibration of model performance across a specific dataset representative of the desired prediction distribution, and utilising the calibration measures to help provide valid error bars for the model output.
@@
In this paper, we conduct a thorough empirical study on conformal prediction for neural-network-based surrogate models to equip their predictions with calibrated, guaranteed error bars, even on out-of-training-distribution scenarios. We demonstrate guaranteed coverage across spatio-temporal domains (high-dimensional outputs), irrespective of the choice of model, data and training conditions. Our work provides a rigorous method to check the usefulness of a pre-trained surrogate model, its validity and applicability in an inference/production scenario.
@@
Pathway to Addressing Climate Change
Machine-learning-based surrogate modelling is key to accelerating scientific simulation as it promises computational efficiency and allows for data-driven discovery at scale. Especially when modelling complex systems, such as computational fluid dynamics, nuclear fusion, and weather forecasting, computational speed of scientific simulation is important; furthermore, it is increasingly crucial to provide accurate, robust, and timely uncertainty estimates. In safety-critical systems Knight (2002), it is impervious that we supplement model predictions with calibrated uncertainty estimations. This can lead to improved decision-making processes in downstream tasks. The CP framework demonstrated in this paper pushes the envelope forward in uncertainty quantification for complex scientific modelling at scale, demonstrating industry-level safety-critical applications.
@@
Nuclear fusion offers a solution to low-carbon, safe, renewable, and cheap energy production. Therefore, solving nuclear fusion at a commercial scale is a key component in decarbonising hard-to-abate industries (Lerede et al., 2023). However, several challenges remain, ranging from confinement(Degrave et al., 2022), and materials (Linke et al., 2019) to mitigating disruptions (Kates-Harbeck et al., 2019). Operating on an accelerated timeline and exponential costs, iterative test-based design becomes an infeasible option. Researchers are left to solve these challenges in-silico using simulations. However, within the context of fusion modelling, at the exascale, simulation costs grow exponentially and need surrogate models capable of quicker, cheaper approximations. Neural-PDE solvers have enabled this, allowing us to solve complex multi-physical systems several orders of magnitude faster than traditional numerical solvers (Pamela et al., 2024; Carey et al., 2024). As further developments are done on these models, it is pertinent to provide rigorous methods of quantifying the uncertainty associated with these models.
Climate change has led to the observation of extreme weather events across the globe (Ebi et al., 2021). Being able to forecast weather rapidly and cheaply allows us to take proactive rather than reactive measures to these climate change-induced events. Neural weather models learnt from both simulation and observed data have emerged as cheaper and faster alternatives to traditional numerical models. However, they suffer from the limitations found within the training data, such as sparse and inaccurate observations as well as the error growth in the numerical models (Sheshadri et al., 2021). Modelling the uncertainty associated with these neural weather predictions is crucial in an operational context (Bouallègue et al., 2024). Through this work, we propose a model-agnostic method that gives calibrated error bounds for all variables, lead times and spatial locations, while not requiring any modifications to the model albeit with negligible computational costs.
@@ 2 Conformal Prediction
Conformal prediction (Vovk et al., 2005; Shafer and Vovk, 2008) answers the following question: Given some arbitrary dataset $(X_{1},Y_{1}),(X_{2},Y_{2}),...,(X_{n},Y_{n})$, and some machine learning model $\hat{f}:\mathcal{X}\to\mathcal{Y}$ trained on this dataset, what is the accuracy of $\hat{f}$ at predicting the next true label $Y_{n+1}$ at query point $X_{n+1}$. CP (Vovk et al., 2005) extends the point prediction $\hat{y}$ of $\hat{f}$ to a prediction set $\mathbb{C}^{\alpha}$, which is guaranteed to contain the true label $Y_{n+1}$ with probability
@@
CP quantifies errors in machine learning since the inequality in Equation 1 is guaranteed regardless of the selected machine learning model and the training dataset $\left\{(X_{i},Y_{i})\right\}_{i=1}^{n}$, other than that the samples used for the calibration (estimating model performance) are exchangeable (a weaker form of i.i.d.).
There have been several variants of CP since its original proposal by (Vovk et al., 2005). Inductive conformal prediction (Papadopoulos, 2008), which we pursue in this work, splits the data into a training set (the usual training set for the underlying ML model) and a calibration set, used to construct prediction set $\mathbb{C}^{\alpha}$ for each $\alpha$ value. The prediction sets are constructed, such that Equation 1 holds, by comparing the trained ML model to the unseen calibration data using a nonconformity score.
@@
Figure 1: Inductive CP framework over a deterministic model (see AER in Section 2.2): (1) Perform calibration using a nonconformity metric (L1 error residual (AER) with $\hat{s}$ representing the calibration scores, $y_{c},\tilde{y}_{c}$ the calibration targets and predictions respectively). (2) Estimate the quantile corresponding to the desired coverage from the CDF of the nonconformity scores ($n$ represents the calibration sample size, $(1-\alpha)$ the desired coverage, $F_{\hat{s}}^{-1}$ the quantile function applied over the inverse CDF of nonconformity scores, $\hat{q}$ the quantile matching the desired coverage). (3) Apply the quantile to the model predictions to estimate the prediction sets ($\tilde{y}_{p}$, the model predictions and $\hat{q}$ the upper and lower bars for the predictions).
The inductive CP framework explored further in this paper follows a three-step procedure as outlined in Figure 1. Initially, we perform a calibration of the model outputs, where a nonconformity metric of the model’s performance is evaluated. This is followed by the quantile estimation, which is performed by obtaining the required quantile from the cumulative distribution of the nonconformity scores obtained in the previous calibration step. The final step involves estimating the prediction sets by applying the estimated quantile over the model predictions in the interested prediction regime.
@@ 2.1.1 Overview
We consider models that describe the evolution of spatio-temporal field variables defined by physical processes, such as those described in numerical PDE modelling, nuclear fusion diagnostics, and weather forecasting. Each modelling task is perceived as an initial value problem, where the calibration and prediction sets are pairs of inputs and outputs characterised by the initial condition and their respective solutions. Within the case of PDE modelling, each datapoint is a single numerical simulation, whereas for experimental modelling (such as nuclear fusion and weather), the datapoint is a single forecast driven by an initial condition.
@@
Within our framework, we perform calibration for each cell individually, estimating the marginal coverage for each cell of the spatio-temporal tensor as demonstrated in Figure 2, i.e., the coverage is guaranteed for each cell taken individually. CP is performed across each spatio-temporal point output by the model, resulting in upper and lower coverage bands for each point. Upon calibrating for the error bars cell-wise, we averaged over each to estimate the coverage across the simulation domain. Within the estimation of the prediction sets, we don’t consider the influence of adjacent field points and implicitly expect the model to extract that within the learning process. We expect the discretised spatio-temporal domain to be the same across the calibration and prediction set.
In this paper, we propose a formulation of the CP framework that provides statistically valid, dimensionality independent marginal coverage over spatio-temporal prediction models. We demonstrate the efficacy of our approach through comprehensive and rigorous experimentation across a range of spatio-temporal models, including complex multi-physics models and state-of-the-art neural weather and nuclear fusion models.
2.1.2 Definition
@@ 2.2 Nonconformity Scores
A nonconformity score can be described as a measure of the model’s performance with respect to the calibration dataset. Thus, the scores are a function of the trained model, the calibration inputs, and outputs (Angelopoulos and Bates, 2023). Nonconformity scores describe the deviation of the model from ground truth. In this paper, we focus on three methods of estimating the nonconformity scores for a model:
@@
  • Quantile Regression (CQR): Outlined by (Romano et al., 2019), three models are trained to perform conformal prediction. The lower and upper coverage bands are estimated by models trained to output the $100\times\alpha^{th}$ and $100\times(1-\alpha)^{th}$ percentile, while the median is learned by training the model to output the $50^{th}$ percentile. During training the models are estimated to output the $\alpha^{th}$ percentile by optimising the quantile loss (Koenker, 2005). The nonconformity score measures the distance of the calibration data point from the nearest coverage band: $s(x,y)=\max\{\underline{f}(x)-y,y-\overline{f}(x)\}$. Using the upper and lower quantiles, CQR offers an initial coverage that can be adjusted and calibrated to perform conformal prediction: Using the required quantile ($\hat{q}$), the prediction set is obtained as $\{\underline{f}(x)-\hat{q},\overline{f}(x)+\hat{q}\}$.
@@
  • Absolute Error Residual (AER): Requiring only a single model, trained to approximate the deterministic function itself, this method uses a more straightforward estimate of the model error to calibrate the coverage. nonconformity scores are estimated by taking the absolute error of the model across the labelled calibration dataset (Lei et al., 2018): $s(x,y)=|y-\tilde{f}(x)|$. Upon deriving the quantile ($\hat{q}$), the prediction set is obtained as : $\{\tilde{f}(x)-\hat{q},\;\tilde{f}(x)+\hat{q}\}$. This method requires no structural changes to the model, and only one deterministic model is required to obtain valid prediction sets. The method uses the distribution of the absolute error residuals (AER) from the model to construct sets with the correct coverage.
@@
  • Standard Deviation (STD): The models are probabilistic, i.e., they predict a distribution rather than a deterministic quantity, where the outputs are characterised by a mean $\mu(x)$ and standard deviation $\sigma(x)$. The nonconformity score, $s(x,y):\frac{|y-\mu(x)|}{\sigma(x)}$, takes into account the uncertainty of the model and conformalises it to provide validity over the predictive uncertainty  [^1] If the choice of the model is not probabilistic in itself, we will have to modify the model with changes to architecture or introduce sampling strategies to output a distribution.. The standard deviation of the output distribution provides estimates of the error bars and; thus, it is an initial measure of the coverage, which can be adjusted to ensure the correct coverage using conformal prediction. Upon deriving the quantile ($\hat{q}$), the prediction set is obtained as : $\{\mu(x)-\hat{q}\sigma(x),\;\mu(x)+\hat{q}\sigma(x)\}$. To do CP using a probabilistic model, we modify our networks to behave in a probabilistic manner through architectural modifications or changing the training regime.
@@
Figure 3 illustrates how these nonconformity scores can help obtain guaranteed coverage, across all modelling scenarios considered in the paper(more details on them in Section 3). The coverage provided across the nonconformity scores may vary, but the CP formulation ensures the validity of the error bars. Irrespective of the choice of the nonconformity score, we obtain near-to-ideal coverage for the prediction following the inequality found within Equation 1. Since coverage is guaranteed in the framework, the choice of the nonconformity score is often undertaken considering other aspects such as architectural modifications, calibration time, and cost of obtaining additional data.